An economy experiencing healthy macroeconomic conditions

a.Is performing below potential; has zero percent overall unemployment rate; and has zero percent cyclical unemployment.

b.Is performing below potential; has unemployment rate equal to the natural rate; and has zero percent frictional unemployment.

c.Is performing at potential; had unemployment rate equal to the natural rate; and has zero percent cyclical unemployment.

d.Is performing at potential; has zero percent overall unemployment rate; and has zero percent cyclical unemployment.

Suppose your uncle started college in 1998 and had a job busing dishes that paid $5 per hour. In 2008 you had the same job; it paid $6 per hour. The CPI in 1998 was 163.0; in 2008 it was 216.5. Which job paid more?

your uncle was paid more in real terms.

you were paid more in real terms.

there is no way of telling who was paid more.

you and your uncle were paid the same in real terms.
Answer #1

Q1:

Under healthy macroeconomic conditions, an economy performs at its potential and has an unemployment rate equal to the natural rate of unemployment.

Natural rate of unemployment = Frictional unemployment + Structural Unemployment

Frictional and Structural types of unemployment are inherent in an economy and can't be completely made zero. Therefore, an economy experiences an unemployment rate equal to its natural rate of unemployment when it performs at its potential.

However, when an economy performs at its potential, the cyclical unemployment becomes zero as the economy is in a boom and all those who have relevant skills and are available to work can find work during such a phase

Ans: c.Is performing at potential; had unemployment rate equal to the natural rate; and has zero per cent cyclical unemployment.

Q2.

Real Wage = (Nominal Wage/CPI in the year) * 100

Nominal Wage in 1998 (your uncle's wage) = $5 per hour

Nominal Wage in 2008 (your wage) = $6 per hour

CPI1998 = 163.0

CPI2008 = 216.5

Real Wage in 1998 = (Nominal Wage in 1998/CPI1998) * 100 = ($5/163) * 100 = $3.067

Real Wage in 2008 = (Nominal Wage in 2008/CPI2008) * 100 = ($6/216.5) * 100 = $2.771

Real wage paid in 2008 is lower compared to 1998. Therefore, you earn lower than your uncle in real terms

Ans: your uncle was paid more in real terms